OpenStudy (anonymous):

If m is an odd integer and n an even integer, which of the following is definitely odd? a..(2m+n)(m-n) b.. (m+n^2)+(m-n^2) c..m^2+mn+n^2 d... m+n

OpenStudy (cwrw238):

is your second sentence correct sauracshakya?

OpenStudy (anonymous):

U know odd number + odd number is even number odd number + even number is odd number odd number times odd number is odd number odd number times even number is even number

OpenStudy (anonymous):

i knew it ....but as you wrote odd +even=even is not correct its odd

OpenStudy (cwrw238):

correct now

OpenStudy (anonymous):

Thanx @cwrw238

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

wt do u think wts the answer

OpenStudy (anonymous):

m=3,n=2 we can substitute to see the answer

OpenStudy (anonymous):

What about d

OpenStudy (anonymous):

there is conflict in between c and d

OpenStudy (anonymous):

Nope a and d

OpenStudy (anonymous):

c is even

OpenStudy (cwrw238):

no a is not odd :- even times odd = even

OpenStudy (anonymous):

Yes not a

OpenStudy (cwrw238):

yea d is definitley odd

OpenStudy (anonymous):

yes a is not odd definately it is even

OpenStudy (anonymous):

im confused about c is it even or odd ...accc to me it is odd

OpenStudy (anonymous):

reply gys

OpenStudy (cwrw238):

m^2 is odd nm is odd and n^2 is even odd + odd + even = even

OpenStudy (anonymous):

nm can not be odd ...take an example 2*3=6 and it is even

OpenStudy (anonymous):

so in this ...c and d both are odd ...,,so its finally done

OpenStudy (cwrw238):

lets look at it again to make sure - i don't trust myself today

OpenStudy (cwrw238):

m^2 is odd mn is even even + odd = odd n^2 is even odd + even = odd yes c is odd

OpenStudy (anonymous):

i was sure about it but in tcs site they did'nt give another option if both answere are correct ...but now i think that there was some mistake

OpenStudy (cwrw238):

yes - thats part of problem you automatically think theres only one solution as u say there could have been mistake in the question.